# Dipper: Diversity in Prompts for Producing Large Language Model Ensembles in Reasoning tasks

### v1

@@ A.1 Prompt Generator
    Fidelity. Each prompt should be able to influence the LLM into applying a certain type of reasoning approach to the task, and not have significant negative impact the LLM’s performance on the task.
@@
    Diversity. The prompts should be sufficiently different from one another such that they elicit various reasoning pathways and provide a diverse pool to select from in the subsequent component.
@@
We first show that LLMs are capable of generating prompts that meet this desideratum, via the most direct way of prompting it to generate a pool of candidate prompts while providing it with exemplars illustrating different reasoning prompts. To do so, we considered a list of 7 reasoning prompts inspired by existing works [25, 26, 27] on prompting methods to boost reasoning capabilities. Given these prompts as exemplars, we used GPT-4o to further generate a set of 200 different candidate prompts that each represent a different reasoning approach (details in Appx. B.1). Fig. 5 shows the distribution of average accuracy over a sampled test set of MATH [24] questions for each prompt, when used for the Qwen2-MATH-1.5B model (i.e., $F(M(\cdot,w);\mathcal{T})$ for $w_{i}\in\mathcal{W}$). Note that the distribution of accuracy is largely higher than that of the base model without prompts, and similar to the accuracies achieved by the reasoning prompt exemplars, demonstrating the fidelity requirement. Qualitatively, we see that the prompts are also relatively diverse – they generally specify certain reasoning approaches inspired by various subject domains (see Table 2). We quantify this diversity in Sec. 3.1 with our proposed metric.
Note that when generating the prompts, we did not pass any task description to the LLM prompt generator. We did so as the reasoning prompts can in general be task-agnostic, even if some may be inspired by some specific subject matter. In practice, the candidate pool of reasoning prompts need not be generated on-the-fly, but be drawn from a shared pool prepared beforehand by a more powerful LLM, to be used by ensembles consisting of much smaller LLMs, as we demonstrated. The actual selection of relevant prompts can be done by the prompt selector component, which we will described next in Sec. 3.1.
@@
Figure 5: The accuracy distribution of 200 candidate prompts on MATH.
Prompt
**Think Like a Lawyer**: Build a case for your answer using evidence and logical arguments.
A.2 Dipper algorithm
Our Dipper algorithm is outlined in Algorithm 1.
Algorithm 1 Dipper semantic volume algorithm
@@ A.2 Dipper algorithm
3:  $\mathcal{Z}\leftarrow\{\ \}$
4:  $\bar{u}(w)\leftarrow[F(M(\cdot,w_{i}),\mathcal{T}_{d})\text{ for }w_{i}\in\bar{\mathcal{W}}]$
5:  $\mathcal{Z}\leftarrow\mathcal{Z}\cup\arg\max_{w}\bar{u}(w)$
6:  $\mathcal{W}\leftarrow\bar{\mathcal{W}}\setminus\arg\max_{w}\bar{u}(w)$
@@
13:        $\tilde{V}_{w_{k}}\leftarrow\log\det(\exp{(\alpha\operatorname*{diag}{u})}RR^{T})$
@@
Appendix B Detailed Experimental Setting
B.1 Fixed 7 prompts and Prompt Generation
@@
B.2 Evaluation
We primarily consider three datasets in our paper. For MATH, we randomly 10% test samples from each category and form a fixed subset of size 500. We uniformly randomly sample 20 samples from this subset to form a validation dataset and use the rest 480 samples as the hold-out test dataset. For GSM8K and MMLU-STEAM, we use their official split of test data and uniformly randomly sample 20 samples to form a validation dataset for each task, and use the rest samples as the hold-out test data.
In the inference evaluation, we use 4-shot exemplars for MATH, 8-shot for GSM8K, and 5-shot for MMLU-STEM. Those exemplars are adopted from the evaluation setting in Qwen2-MATH [28] and fixed for all questions and all methods.
B.3 LLM aggregation
We perform LLM aggregation with the same Qwen2-MATH-1.5B-it model, by feeding the question context and the responses from LLM agents into the designed template shown below in the bounding box:
Table 5: The prompt template for LLM aggregation.
[…]
The question is:
[…]
There are some responses:
[…]
Examine those responses and provide the final answer.

### v2

@@ 4.2 Prompt Generator
    Fidelity. Each prompt should be able to influence the LLM into applying a certain type of reasoning approach to the task without significantly degrading the task performance.
@@
    Diversity. The prompts should differ sufficiently to elicit various reasoning pathways and provide a diverse selection pool for the subsequent component.
@@
We first show that LLMs are capable of generating prompts that meet these desiderata, via the most direct way of prompting it to generate a pool of candidate prompts while providing it with exemplars illustrating different reasoning prompts. To do so, we considered a list of 7 reasoning prompts inspired by existing works (Wang et al., 2023a; Deng et al., 2023; Yao et al., 2022) on prompting methods to boost reasoning capabilities. Given these prompts as exemplars, we used GPT-4o to further generate a set of 200 different candidate prompts that each represent a different reasoning approach (details in Appx. A.1).
@@
Figure 1: The accuracy distribution of 200 candidate prompts on MATH with Qwen2-MATH-1.5B.
Fig.˜1 shows the distribution of average accuracy over a sampled test set of MATH (Hendrycks et al., 2021) questions for each prompt, when used with the Qwen2-MATH-1.5B model (i.e., $F(M(\cdot,w);\mathcal{T})$ for $w_{i}\in\mathcal{W}$). Note that the distribution of accuracy is largely higher than that of the base model without prompts, and similar to the accuracies achieved by the reasoning prompt exemplars, demonstrating the fidelity requirement. Qualitatively, we see that the prompts are also relatively diverse – they generally specify certain reasoning approaches inspired by various subject domains (see Appx. C.6). We will quantify this diversity in Sec. 4.3 with our proposed metric.
Note that when generating the prompts, we did not pass any task description to the LLM prompt generator. We did so as the reasoning prompts can be task-agnostic. In practice, the candidate pool of reasoning prompts need not be generated on-the-fly, but can be drawn from a shared pool prepared beforehand by a more powerful LLM, to be used by ensembles consisting of much smaller LLMs, as we demonstrated. The actual selection of relevant prompts from this larger pool can then be done by the prompt selector component, which we will describe next in Sec. 4.3.
4.3 Prompt Selector
With our framework, the optimization problem in Eq.˜2 reduces to an optimization to choose the best subset of prompts $\{w_{i}\}_{i=1}^{n}$ from the set of candidate prompts $\mathcal{W}$:
$\operatorname*{arg\,max}_{\{w_{i}\in\mathcal{W}\}_{i=1}^{n}}F(\mathcal{E}(q_{t};M,n,\{w_{i}\}_{i=1}^{n}),\mathcal{T}).$ (3)
Unfortunately, directly optimizing Eq.˜2 is a combinatorial problem that is very challenging, even if a development/validation set is available for the task of interest. For example, selecting 5 prompts from a candidate pool of 200 prompts involves searching over ${200\choose 5}\approx 2.5\times 10^{9}$ candidates. Instead, we note that the best ensemble composition requires a balance of the two desiderata: fidelity and diversity. Hence, we propose optimizing Eq.˜2 by considering how to prioritize the prompts that have the best predicted performance on the task $\mathcal{T}$, while maximizing the diversity of the selected set of prompts. Our method draws inspiration from past works on determinantal point processes (DPP) (Kulesza, 2012; Lau et al., 2025a), which consider similarity kernels comprising separate quality and diversity terms that match our requirements.
Prompt fidelity.
First, we can approximate the predicted performance of each prompt by its average performance on a task development set $\mathcal{T}_{d}$ [^1] Without such a development set, an uninformed prior on the performance (e.g. uniform distribution across roles), or an informed-prior based on domain knowledge, could also be used.. Note that as inference using these various prompts on a small development set can be done in parallel, this process can in practice be significantly sped up by existing batch inference techniques such as those employed by vLLM (Kwon et al., 2023).
Specifically, for a candidate pool of prompts $\mathcal{W}$ and development set $\mathcal{T}_{d}$, we can define a prompt fidelity mapping $u:\mathcal{W}\rightarrow[0,1]$,
$u(w)\coloneq F(M(\cdot,w),\mathcal{T}_{d}),$ (4)
where $M(\cdot,w)$ is the LLM model conditioned by prompt $w\in\mathcal{W}$, and $F$ the expected accuracy defined in Section˜3. In practice, for a candidate pool of size $n$, $u(w)$ can be represented as an $n\times 1$ column vector, with the elements representing each prompt’s expected accuracy.
Semantic entropy.
Then, we measure prompt diversity by considering how different the semantic meanings of the $n$ role prompts are from each other. We represent each prompt’s semantic meaning with a mapping $R$ from its text representation $w$ into a normalized continuous vector $s\in\mathbb{R}^{p}$ in a $p$-dimensional semantic embedding space $\mathcal{S}$ through a sentence embedding model $M_{s}$ (Reimers and Gurevych, 2019), i.e., $R(w)\coloneq M_{s}(w)$. This mapping can be represented as an $n\times p$ prompt embedding matrix $R=[s_{1},\cdots,s_{n}]$ where $s$ is a $1\times p$ row vector representing each prompt.
To quantify prompt diversity of a given set of prompts, we propose to compute the volume enclosed by the selected prompts in semantic space. Intuitively, for $n$ fixed prompts, more diverse prompts point to more varied directions in semantic space, and enclose a larger volume. Specifically, we note that from basic geometry, the determinant of a Gram matrix is the squared volume of the parallelepiped spanned by the embedding vectors. Hence, we define the semantic volume metric $V$ as
$V\coloneq\log\det(RR^{T}),$ (5)
where we take the logarithm (for numerical stability) of the Gram matrix determinant [^2] We omit a factor of 2 which does not affect the optimization process. For our setting, we also have $n<p$ as the semantic embedding space is usually high-dimensional. Sec. C.5 shows how sets of prompts that are qualitatively observed to be more diverse have larger quantitative semantic volume.
Fidelity-adjusted semantic volume (FASV).
To incorporate the prompts’ expected accuracy information, we can compute the performance-adjusted prompt embedding matrix,
$\tilde{R}\coloneq\exp({\frac{\alpha}{2}\operatorname*{diag}(u)})R,$ (6)
where $\operatorname*{diag}(u)$ is the diagonal matrix with its $i^{\text{th}}$ diagonal element being the corresponding element $u_{i}$. This essentially scales each row $s_{i}$ in $R$ by an exponential factor based on its corresponding predicted accuracy, $\exp({\frac{\alpha}{2}u_{i}})$, where $\alpha$ is a scalar hyperparameter influencing the balance between diversity and expected performance. Intuitively, prompts with higher expected accuracy would then be able to support larger semantic volume and hence be prioritized for inclusion into the ensemble. The adjusted embedding matrix can then be used to compute the semantic volume in Eq.˜5, which simplifies to
$\tilde{V}=\log\det(\tilde{R}\tilde{R}^{T})=V+\alpha\|u\|_{1}\ ,$ (7)
providing an interpretable expression illustrating the balance between the diversity (i.e., the semantic volume metric in Eq.˜5) and fidelity desiderata (i.e., the L1 norm of the prompt fidelity metric in (Eq.˜4) that needs to be optimized for the ensemble. Derivation details are in Appendix˜B, and we provide empirical analysis of the effectiveness of this combined metric in Section˜5.3.
Algorithm 1 Dipper FASV algorithm
@@ 4.3 Prompt Selector
[…]
@@
13:   $\tilde{V}_{w_{k}}\leftarrow\log\det(RR^{T})+\alpha\|u\|_{1}$
@@
Optimization of semantic entropy.
We can now recast Eq.˜2 as an optimization of the fidelity-adjusted semantic volume metric $\tilde{V}$ evaluated over the set of candidate prompts. Note that instead of the expected ensemble performance $F(\mathcal{E})$, which is an objective that can only be optimized by blackbox optimization methods like Bayesian Optimization (Garnett, 2023; Dai et al., 2023; Lau et al., 2025b), our metric $\tilde{V}$ can be efficiently approximated by well-established heuristics.
Specifically, as the semantic volume metric is submodular, we can optimize for the best subset of roles by incrementally building the subset with a greedy approach up to the desired size $n$ and still be guaranteed a good approximation (Nemhauser et al., 1978). This is an important advantage that allows us an efficient and theoretically-inspired approach to obtain the best ensemble prompts. Our proposed algorithm is outlined in Algorithm˜1.
4.4 Response Aggregator
Given the responses from the various LLMs of constituents, the aggregation method determines how much information from the constituents is used to derive the final output of the ensemble. We consider the two most popular approaches:
Majority voting (MV)
It involves extracting the final answer $\hat{c}$ from each LLM response $\hat{y}=\{\hat{r},\hat{c}\}$, and selecting the answer that has been proposed the most number of times. This approach does not evaluate the quality of reasoning $\hat{r}$ output produced by each LLM, but is easily implementable.
Best-of-N
An external reward model is implemented to evaluate the response $\hat{y}$ of each agent, and the response with the highest score is selected as the final response. This approach does not leverage consensus among constituents but could be effective in identifying the correct responses that would be only covered by a few agents.
5 Experiments
Experimental set-up. We empirically evaluate our framework on mathematically reasoning tasks with the MATH (Hendrycks et al., 2021), GSM8K, and MMLU-STEM datasets. We implement our framework using the GPT-4o as our prompt generator and Qwen2-MATH-1.5B as the constituent model in the ensemble, where the ensemble constituents are run in parallel using vLLM (Kwon et al., 2023) for fast batch inference. Further details (Appx. A) and additional results (Appx. C) are in the Appendix.
Baselines.
We evaluate our Dipper framework by comparing it against the "Self-ensemble" method, which lacks prompt diversity but incorporates diversity through repeated response sampling (Wang et al., 2023b), along with the single model performance as a reference. We also include two other implementation variants of Dipper in our analysis, beyond the implementation based on semantic volume, "Dipper (FASV)":
  1. 1.
    Random+. Here we randomly sample prompts from the candidate pool based on a probability distribution proportional to their predicted accuracy as defined in Eq.˜4, i.e., $p(w)\propto u(w)$. This aims to achieve diversity through the sampling process while prioritizing prompts with higher predicted accuracy.
  2. 2.
    Top-n. Here we greedily select the top $n$ prompts which are ranked based on their predicted accuracy $u(w)$. It assumes that the diversity of prompts introduced by our prompt generation process is sufficient and hence does not explicitly optimize for ensemble diversity during the prompt selection phase.
Figure 3: Accuracy vs. average number of unique answers using different numbers of prompts in ensembles.
5.1 Ensembles with fixed prompt methods
To motivate our Dipper framework and demonstrate the importance of prompt diversity, we first consider a fixed set of seven distinct reasoning prompts inspired by existing works (Wang et al., 2023a; Deng et al., 2023; Yao et al., 2022) (details in Appx. A.1). With a fixed ensemble size of seven, Fig.˜2 shows that an ensemble using these seven different prompts (57.31%) outperforms both a baseline self-ensemble without prompt variation (55.76%) and the average performance (56.55%) of seven self-ensembles, each using only one of the distinct prompts.
In addition, we evaluated the impact of prompt diversity by constructing ensembles with varying numbers of unique prompts (from one to seven) drawn from this set, while maintaining an ensemble size of seven. When fewer than seven unique prompts were used, responses were randomly sampled to meet the ensemble size. The result in Fig.˜3 indicates that increasing the number of unique prompts generally leads to higher accuracy and reduced variance. This suggests that prompt diversity within an ensemble can enhance performance and consistency, particularly when the performances of prompts are unknown before the final evaluation.
Figure 4: Comparison of different ensemble methods on MATH for Qwen2-MATH-1.5B.
Figure 5: Comparison of different ensemble methods on MATH for the LLaMA3.2B model.
5.2 Ensembles with optimized prompt diversity
Next, we consider our full Dipper framework. We first generate a pool of prompt candidates ($|\mathcal{W}|=200$) using the 7 reasoning prompts in the previous section as in-context exemplars (details in Appx. A.1) and then perform prompt fidelity-diversity optimization (Sec. 4.3) to select the best ensemble prompts. As shown in Fig.˜5, our full Dipper implementation with FASV achieves the highest accuracy compared to the self-ensemble baseline and all other Dipper variants across various ensemble sizes. Dipper also significantly outperforms the single LLM. For example, Dipper with $n=9$ has close to a 10%-pt increase (~20% accuracy gain) compared to the single LLM baseline. In fact, our ensemble that consists of just 3 Qwen2-MATH-1.5B models already (slightly) outperforms the next model size class, the Qwen2-MATH-7B model. Note also that the performance gain of Dipper over the self-consistency baseline is about as large as the gain from moving up one model class (from the 1.5B to 7B model). We see similar results on MATH with the general model LLaMA3.2B in Fig. 5, where Dipper (FASV) is shown consistently effective. Refer to Appendix˜C for more results for other datasets (e.g., GSM8K, MMLU-STEM, BIG-Bench) and models.
5.3 Fidelity-diversity optimization
To further understand the mechanisms behind Dipper’s performance gains, we analyze the predictive power of our fidelity-adjusted semantic volume metric in Eq.˜7 (which we denote as $V$ in this section for notational simplicity) on the final ensemble performance on the test set $F(\mathcal{E})$. We quantify this by computing the Spearman correlation between $V$ and $F(\mathcal{E})$: the higher the Spearman correlation, the better our optimization of the ensemble prompts via $V$ will lead to higher ensemble test performance. Fig.˜6 shows the Spearman correlation of $V$ and $F(\mathcal{E})$ for the MATH dataset experiment, with different fidelity-diversity hyperparameter $\alpha$ values. We can observe two key insights.
Figure 6: Spearman correlation between $V$ and test performance $F(\mathcal{E})$ on the MATH under different fidelity-diversity hyperparameter $\alpha$.
Figure 7: The scatter plot showing correlation between $V$ and $F(\mathcal{E})$ for different prompt candidates.
First, there is a relatively strong positive correlation between $V$ and $F(\mathcal{E})$, going as high as $0.8$ for some values of $\alpha$. This corroborates our main results where our Dipper method that explicitly optimizes for $V$ outperforms other baselines and achieves higher $F(\mathcal{E})$.
Second, there is a U-shape trend between the Spearman correlation and hyperparameter value $\alpha$, where the correlation increases as $\alpha$ increases from 0, but decreases after a certain point. This trend demonstrates the need of taking into account both fidelity and diversity when optimizing for the set of ensemble prompts, as we discussed in Section˜4.3. On the one hand, $\alpha=0$ corresponds to the case where we focus solely on diversity and ignore the fidelity or individual predicted performance of prompts ($u(w)$) – this may select a set of diverse prompts, but potentially some irrelevant or poor performing prompts. However, if we emphasize fidelity too much and disregard diversity, we may end up selecting very similar prompts resulting in less ensemble performance gains. At the extreme, choosing large $\alpha$ reduces to the Top-n baseline implementation, which has poorer performance than Dipper that optimizes for semantic volume. In practice, just like other machine learning hyperparameters, we could inform the choice of $\alpha$ with the development set, if available.
5.4 Prompt candidate matters
We then analyze how the candidate pool diversity introduced by Prompt Generator contributes to our framework. Out of the original candidate set $\mathcal{W}$ (Candidate set 2), we obtained another set by selecting one cluster after performing k-means clustering over $\mathcal{W}$ with $k=4$ ($\mathcal{W}^{\prime}$, Candidate set 1). We then randomly select ensembles of size $n=5$, and plot their respective $V$ and $F(\mathcal{E})$ in Fig.˜7. We can see that ensembles from $\mathcal{W}^{\prime}$ have much lower accuracy and semantic volume compared to those from $\mathcal{W}$, illustrating the importance of the candidate pool diversity from the Prompt Generator.
5.5 Dipper combined with other prompting methods like Reflexion
In addition, we also show that our ensemble framework Dipper is orthogonal to other established prompting techniques (e.g. CoT and Reflexion (Shinn et al., 2024)), allowing it to stack and bring greater performance. To demonstrate this, we first use Dipper to select 5 agents and query each agent with questions from the MATH dataset. Their initial responses will then be self-reflected according to the method proposed in Reflexion (Shinn et al., 2024), before being aggregated into the final answer with MV. We found that combining self-reflection with Dipper achieves a performance gain of $8\%$ (from an accuracy of $57\%$ to $65\%$), demonstrating that Dipper has the potential to be extended further or combined with other methods.
Figure 8: Comparison of Dipper and self-ensemble baseline on MATH using LLaMA-3B and Best-of-N aggregation.
5.6 Generalization to Best-of-N aggregation
Finally, we study the effects of using the Best-of-N aggregation for our response aggregator component, showing that Dipper can work well with external reward models. We use an existing reward model, “Qwen2.5-Math-RM-72B” to assess the quality of the generated responses and select the final response, using the LLaMA-3B model and Best-of-N aggregation in our Dipper framework. As seen in Fig.˜8, the Dipper variants significantly beats the self-ensemble baseline when Best-of-N is used. In this scenario, the performance among the Dipper variants are closer since Best-of-N considers each constituent individually rather than jointly (like in MV) to produce the final answer, though our full Dipper implementation still consistently performs the best. We also see that Dipper can stack with benefits from a strong verifier model, given its performance gains compared to the result in Fig.˜10 where no verifier model is available.
6 Conclusion
In this work, we have proposed a novel framework, Dipper, where a single LLM model type is fed an optimized, diverse set of reasoning prompts in parallel, effectively producing an ensemble at inference time to achieve performance improvement in reasoning tasks. Our empirical findings have demonstrated the effectiveness of various Dipper implementations in improving inference performance for a variety of reasoning tasks, which may inspire future works to investigate additional optimization methods for prompt-based inference-time ensembles to further improve performance gains.
Limitations
Our framework Dipper focuses on developing inference-time ensembles where each constituent is based on the same base model – this caters to the most common and straightforward scenario where users are using a single LLM model and can apply Dipper to further boost its reasoning performance at inference time without additional training. However, when users may wish to use heterogeneous models, Dipper currently does not take into account such model diversity, which we believe may enable further performance boosts if properly optimized. We leave it to future works to potentially build on Dipper to extend it beyond its current limitations in this regard.
Acknowledgment
This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD/2023-01- 039J). This research is part of the programme DesCartes and is supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. This research/project is supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative (AISG Award No: AISG-NMLP-2024-001). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore. This research/project is also supported by SAP and Singapore’s Economic Development Board under the Industrial Postgraduate Programme.
Appendix A Detailed Experimental Setting
The huggingface model path for the primary model we used is “Qwen/Qwen2-Math-1.5B-Instruct” and the sentence transformer’s model path is ‘all-MiniLM-L6-v2”. We use the default generation parameters for Qwen2-Math-1.5B-Instrct: the temperature is set to 0.7, the top probability used for filtering is set to 0.8, and the repetition penalty is set to 1.05. We also set the max tokens to be generated to 512.
A.1 Fixed 7 prompts and Prompt Generation
@@
A.2 Evaluation
We primarily consider three datasets in our paper. For MATH, we randomly sample 10% test samples from each category in its official test split and form a fixed subset of size 500. We then uniformly randomly sample 20 samples from this subset to create a validation dataset and use the rest 480 samples as the hold-out test dataset. For GSM8K and MMLU-STEM, we use their official split of test data and uniformly randomly sample 20 samples to form a validation dataset for each task, and use the rest samples as the hold-out test data.
In the inference evaluation, we use 4-shot exemplars for MATH, 8-shot for GSM8K, and 5-shot for MMLU-STEM. Those exemplars are adopted from the evaluation setting in Qwen2-MATH (Team, 2024) and fixed for all questions and all methods.
Appendix B Fidelity-adjusted semantic volume metric
In this section, we provide the explicit derivation of how our fidelity-adjusted semantic volume metric can be simplified to a weighted sum of two terms representing the diversity and fidelity desiderata in Eq.˜7, which clearly illustrates the balance between the two desiderata during the optimization process.
$\tilde{V}$= $\log\det(\tilde{R}\tilde{R}^{T})$ (8)
$=$ $\log\det\left(\exp({\frac{\alpha}{2}\operatorname*{diag}(u)})R\right)\left(\exp({\frac{\alpha}{2}\operatorname*{diag}(u)})R\right)^{T}$ (9)
$=$ $\log\left[\det\left(\exp({\frac{\alpha}{2}\operatorname*{diag}(u)})\right)\det\left(RR^{T}\right)\det\left(\exp({\frac{\alpha}{2}\operatorname*{diag}(u)})^{T}\right)\right]$ (10)
$=$ $\log\det(RR^{T})+2\log\det\left(\exp({\frac{\alpha}{2}\operatorname*{diag}(u)}\right)$ (11)
$=$ $V+2\log\prod_{i}\exp({\frac{\alpha}{2}u_{i}})$ (12)
$=$ $V+\alpha\|u\|_{1}\$(13)
where Eq.˜9 follows from the definition of $\tilde{R}$ in Eq.˜6, Eq.˜10 uses the identity $\det(AB)=\det(A)\det(B)$, Eq.˜11 the identity $\log(AB)=\log(A)+\log(B)$, Eq.˜12 the definition of semantic volume in Eq.˜5, and Eq.˜13 noting that $\sum_{i}u_{i}=\|u\|_{1}$ since $u\geq 0$.